How much force is needed to accelerate a 66kg skier 1 m/s2.
1
Expert's answer
2020-11-23T10:26:18-0500

By the definition of the Newton's Second Law of Motion, we have:


"F=ma=66\ kg\cdot 1\ \dfrac{m}{s^2}=66\ N."

Answer:

"F=66\ N."